# Quantitative imaging of nonlinear spin-wave propagation using diamond quantum sensors

### v1

@@ III.6 Comparison with a different YIG film
So far, we have quantitatively imaged the amplitude and wavenumber distribution of spin waves in YIG-A over a wide range of microwave input powers. In particular, we have confirmed that the threshold of nonlinearity coincides with the threshold determined by four-magnon scattering. As shown in Eq. (14), this threshold depends on the relaxation rate of the spin waves. It is vital to investigate how our interpretation can be generalized. Therefore, we present the results of YIG-B with a different relaxation rate.
@@
The external magnetic field $B_{0}$ is set to $22.45\,\mathrm{mT}$ when measuring YIG-B. The resonance frequency of the NV spins $f_{\mathrm{NV}}$ is $2243.1\,\mathrm{MHz}$, and the wavenumber of the surface spin waves at this frequency is around $1.3\,\mathrm{rad/\mu m}$ (see Supplemental Materials for details).
@@
We sweep the input microwave power $P_{\mathrm{mw}}$ through 23 points from 00 dBm to $22$ dBm. Figure 5(a) displays the microwave amplitude distribution $B_{\mathrm{mw}}(x)$ for five representative input powers (see Supplemental Materials for additional data). The triangular markers with error bars indicate the experimental data. The integrated region along the $y$-direction is the same as that used for YIG-A [see Figs. 1(d) and 2(a)]. Focusing on the data at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$, we observe that, compared to YIG-A, the decay of $B_{\mathrm{mw}}(x)$ from the left to the right end is small, reflecting a lower relaxation rate of the spin waves. Overall, the mean line of $B_{\mathrm{mw}}(x)$ attenuates by approximately 20% over the entire field of view, which corresponds to a decay length of about $300\,\mathrm{\mu m}$.
@@
Figure 5: (a) Input microwave power dependence of the one-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ obtained for YIG-B. The triangular markers with error bars represent experimental data, and the solid black lines correspond to the fitting results. (b) Comparison of $B_{\mathrm{mw}}(x)$ for data with large input microwave power ($P_{\mathrm{mw}}=22\,\mathrm{dBm}$) and data with small input microwave power ($P_{\mathrm{mw}}=6\,\mathrm{dBm}$) for YIG-B. (c) Microwave amplitude distribution and (d) Wavenumber distribution from the spin waves for each microwave input power calculated from the fitting results for YIG-B.
@@
In Fig. 5(b), $B_{\mathrm{mw}}(x)$ at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$ and $P_{\mathrm{mw}}=22\,\mathrm{dBm}$ are overlaid for direct comparison [see Fig. 2(b) for YIG-A]. Although there is no large decay at $P_{\mathrm{mw}}=22\,\mathrm{dBm}$, the amplitude of the mean line is reduced by approximately 50% from the left to the right end of the field of view, indicating the emergence of nonlinear decay at this input power. Furthermore, when focusing on the smaller $x$-coordinate shown in the inset, a slight difference in the wavenumber between the two microwave input powers is observed. However, the modulation is less significant than that seen in YIG-A.
@@
We follow the same procedure as YIG-A to fit $B_{\mathrm{mw}}(x)$, with two differences. First, for YIG-B, across all regions and at every microwave input power, the reference microwave distribution is consistently smaller than the measured microwave amplitude distribution (see Supplemental Materials for details). Consequently, we employ Eq. (3) for the fitting. Second, because the observed wavenumber modulation is small, we adopt the following linear approximation for the wavenumber distribution to describe the minor spatial modulation:
@@
For each microwave input power, the fitting result reproduces the experimental data precisely, as shown by black solid lines in Fig. 5(a).
@@
In Fig. 5(c), we plot the microwave amplitude distribution from spin waves $B_{\mathrm{sw}}(x)$, estimated from the fitting results at eleven different microwave input powers. The left end of the plot is set at $x=0.91\,\mathrm{\mu m}$. Notably, the amplitude at the left end continues to increase up to the maximum input power of $P_{\mathrm{mw}}=22\,\mathrm{dBm}$. In contrast, the amplitude at the right end saturates at around $B_{\mathrm{sw}}=0.5\,\mathrm{mT}$, indicating the onset of nonlinearity. Figure 5(d) shows the estimated wavenumber distribution $k_{x}(x)$ for each input microwave power. The modulation of $k_{x}(x)$ across the field of view is minimal; even at the maximum microwave input power $P_{\mathrm{mw}}=22\,\mathrm{dBm}$, the variation from the left to the right end is less than $0.1\,\mathrm{rad/\mu m}$. These observations are very different from those in YIG-A, as shown in Figs. 2(c) and (d).
@@
Figure 6: (a) Calculated spin-wave amplitude distribution $m_{x}(x)$ for each microwave input power for YIG-B. (b) The heatmap of the spin-wave amplitude distribution for each microwave input power for YIG-B. (c) (d) Dependence of the spin-wave amplitude on input microwave amplitude (c) at $x=4\,\mathrm{\mu m}$ and (d) at $x=78\,\mathrm{\mu m}$ for YIG-B.
@@
As we did for YIG-A, we use Eq. (5) to convert $B_{\mathrm{sw}}(x)$, shown in Fig. 5(c), into the corresponding spin-wave amplitude distributions $m_{x}(x)$. Figure 6(a) shows the converted spin-wave amplitude, and Fig. 6(b) displays a heatmap of $m_{x}(x)$ for all data, along with the individual profiles from Fig. 6(a). To investigate the nonlinearity of the spin waves, we examine the dependence of the spin-wave amplitude on the microwave input amplitude at two positions: $x=4\,\mathrm{\mu m}$ on the left side of the field of view and $x=78\,\mathrm{\mu m}$ on the right side. In Fig. 6(c), the dependence of the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ on the microwave input amplitude is shown. The red dashed line represents a linear fit through the origin, obtained using data from the five smallest microwave input amplitudes (up to $P_{\mathrm{mw}}=10\,\mathrm{dBm}$). The data indicate that nonlinearity begins to occur when the spin-wave amplitude exceeds approximately $10\,\mathrm{mT}$. Although the amplitude does not immediately saturate, it deviates from the linear trend and the gradient continues to decrease. Similarly, Fig. 6(d) shows the dependence of the spin-wave amplitude at $x=78\,\mathrm{\mu m}$. As with the data at $x=4\,\mathrm{\mu m}$, nonlinearity appears when the amplitude reaches around $10\,\mathrm{mT}$. However, in contrast to the left side, the amplitude at $x=78\,\mathrm{\mu m}$ does not increase much in the nonlinear region and appears to saturate around $13\,\mathrm{mT}$. These behaviors are qualitatively similar to those observed in YIG-A, but the threshold at which nonlinearity emerges is smaller than in YIG-A.
@@
We compare the obtained results with the theoretical threshold given by Eq. (14). The right column of Table 1 lists the parameters used in the calculation. The saturation magnetization is taken as $M_{\mathrm{s}}=162\,\mathrm{mT}$, as measured by SQUID magnetometry. The anisotropy field is estimated from the difference between the effective magnetization $M_{\mathrm{eff}}=177.6\,\mathrm{mT}$ and the saturation magnetization. The decay rate is calculated based on the decay length $l_{d}=318\,\mathrm{\mu m}$, obtained from the fitting result at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$. With these parameters, the threshold spin-wave amplitude for the onset of four-magnon decay is estimated to be $\mu_{0}m^{\mathrm{th}}_{x}=10.5\,\mathrm{mT}$. As expected, this threshold is lower than that of YIG-A, reflecting the smaller decay rate in the present case. As shown in Figs. 6(c) and (d), the experimental data indicate that nonlinearity begins to occur when the spin-wave amplitude exceeds approximately $10\,\mathrm{mT}$ (marked by the horizontal lines in the graphs), again in good agreement with this theoretical prediction.
@@ IV Conclusion and Perspectives
To conclude, we employed QDM to investigate the nonlinear dynamics of surface spin waves quantitatively. For YIG-A, we swept the microwave input power to map the microwave amplitude distribution, revealing both nonlinear behavior in the amplitude profile and shifts in the wavenumber resulting from a reduction in the saturation magnetization. Quantitative analysis of these data showed that nonlinearity in the amplitude emerges once the spin-wave amplitude exceeds a specific threshold, consistent with theoretical predictions based on four-magnon scattering. Furthermore, as the amplitude of the spin waves increases, we observed the modulation of the wavenumber within the field of view, and our quantitative analysis revealed that this modulation largely derives from the spin waves generated by multi-magnon scattering. Similar measurements conducted on YIG-B, which exhibits a different relaxation rate, confirmed that the threshold for the onset of nonlinearity also agrees with theoretical predictions.
@@
These results deepen our understanding of nonlinear spin-wave dynamics and underscore the utility of NV centers as sensors for the quantitative investigation of nonlinear spin-wave phenomena, which were previously limited to qualitative assessments.
@@
As a future extension of this work, one can consider the quantitative visualization of the incoherent magnons generated by four-magnon scattering. NV centers can quantitatively measure the magnetic noise at their resonance frequency via longitudinal relaxation measurements. Therefore, by exciting surface spin waves with frequencies detuned from the NV spin resonance and monitoring the relaxation rate of the NV spins while varying the input microwave power, it becomes possible to extract quantitative information about the incoherent magnons produced by nonlinear scattering [35]. Also, at the same condition, quantitative detection of the coherent spin waves detuned from the NV resonance frequency is possible by employing the sensing based on the AC Zeeman effect [26].
@@
In addition, nonlinear phenomena regarding spin waves are not limited to surface spin waves. For example, parametric excitation, where spin waves are generated by microwave fields at twice the frequency of the spin waves oriented along the saturation magnetization direction [3, 4, 5], is an interesting phenomenon. Furthermore, in the case of large-amplitude spin waves, the propagation of spin-wave bullets, which behave as solitons [38], represents another example. It is anticipated that NV centers can also be effectively employed to investigate these phenomena. The nonlinear dynamics of spin waves are also anticipated to have applications in spintronics devices such as transistors [39] and phase shifters [40]. Methods capable of quantitatively evaluating these operations can also provide valuable guidelines for device design as well.
@@ Acknowledgements
This work was partially supported by JST, CREST Grant No. JPMJCR23I2, Japan; Grants-in-Aid for Scientific Research (Nos. JP22K03524, JP23H01103, JP22KJ1058, JP22KJ1059, and JP24H01666); the Mitsubishi Foundation (Grant No. 202310021); Seiko Instruments; Advanced Technology Foundation Research Grants; the Cooperative Research Project of RIEC, Tohoku University; “Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM)” (No. JPMXP1222UT1131) of the Ministry of Education, Culture, Sports, Science and Technology of Japan (MEXT). K.O. and M.T. acknowledge supports from FoPM, WINGS Program, The University of Tokyo.

### v2

@@ III.6 Comparison with a different YIG film
So far, we have quantitatively imaged the amplitude and wavenumber distributions of spin waves in YIG-A over a wide range of input microwave powers. In particular, we have confirmed that the threshold of nonlinearity coincides with the threshold determined by four-magnon scattering. As shown in Eq. (14), this threshold depends on the relaxation rate of the spin waves. It is therefore important to investigate how our interpretation can be generalized. One possible approach to demonstrate the validity of our results is to check whether the amplitude indirectly inferred in Ref. [13] is consistent with theory. However, a straightforward and quantitative comparison is difficult for two reasons. First, the present work investigates a two-dimensional extended YIG thin film, whereas Ref. [13] studies a waveguide. As a result, key parameters of the magnon Hamiltonian, including dipolar interactions, differ substantially, making it difficult to apply the theory used in our paper as is. Second, in Ref. [13], the spin-wave precession angle is inferred from a numerically computed dispersion by imposing a uniform tilt $\theta$ and then allowing the spins to precess. In this procedure, however, the ellipticity of spin-wave precession is not taken into account, and it is unclear how faithfully the imposed $\theta$ maps onto the actual dynamic magnetization amplitude in the experimental system. For these reasons, the indirectly determined threshold reported in Ref. [13] cannot be assessed in a rigorous manner. Instead, to substantiate the validity of both the theory and the measurements, we present results of YIG-B, which has a different relaxation rate.
@@
The external magnetic field $B_{0}$ is set to $22.45\,\mathrm{mT}$ when measuring YIG-B. The resonance frequency of the NV spins $f_{\mathrm{NV}}$ is $2243.1\,\mathrm{MHz}$, and the wavenumber of the surface spin waves at this frequency is around $1.3\,\mathrm{rad/\mu m}$ (see Supplemental Materials for details [30]).
@@
We sweep the input microwave power $P_{\mathrm{mw}}$ in 23 steps from $0\,\mathrm{dBm}$ to $22\,\mathrm{dBm}$. Figure 5(a) displays the microwave amplitude distribution $B_{\mathrm{mw}}(x)$ for five representative input powers (see Supplemental Materials for additional data [30]). The triangular markers with error bars indicate the experimental data. The integration region along the $y$-direction is the same as that used for YIG-A [see Figs. 1(d) and 2(a)]. Focusing on the data at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$, we observe that, compared with YIG-A, the decay of $B_{\mathrm{mw}}(x)$ from the left to the right end is small, reflecting a lower relaxation rate of the spin waves. Overall, the mean line of $B_{\mathrm{mw}}(x)$ attenuates by approximately 20% across the entire field of view, which corresponds to a decay length of about $300\,\mathrm{\mu m}$.
@@
Figure 5: (a) Input microwave power dependence of the one-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ obtained for YIG-B. The triangular markers with error bars represent experimental data, and the solid black lines correspond to the fitting results. (b) Comparison of $B_{\mathrm{mw}}(x)$ for data with large input microwave power ($P_{\mathrm{mw}}=22\,\mathrm{dBm}$) and data with small input microwave power ($P_{\mathrm{mw}}=6\,\mathrm{dBm}$) for YIG-B. (c) Microwave amplitude and (d) Wavenumber distribution from the spin waves for each input microwave power calculated from the fitting results for YIG-B.
@@
In Fig. 5(b), $B_{\mathrm{mw}}(x)$ at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$ and $P_{\mathrm{mw}}=22\,\mathrm{dBm}$ are overlaid for direct comparison [see Fig. 2(b) for YIG-A]. Although there is no pronounced decay at $P_{\mathrm{mw}}=22\,\mathrm{dBm}$, the amplitude of the mean line is reduced by approximately 50% from the left to the right end of the field of view, indicating the emergence of nonlinear decay at this input power. Furthermore, when focusing on the smaller $x$-coordinate shown in the inset, a slight difference in wavenumber between the two input microwave powers is observed. However, the modulation is less significant than that seen in YIG-A.
@@
We follow the same procedure as YIG-A to fit $B_{\mathrm{mw}}(x)$, with two differences. First, for YIG-B, across all regions and at every input microwave power, the reference microwave distribution is consistently smaller than the measured microwave amplitude distribution (see Supplemental Materials for details [30]). Consequently, we employ Eq. (3) for the fitting. Second, because the observed wavenumber modulation is small, we adopt the following linear approximation for the wavenumber distribution to describe the minor spatial modulation:
@@
For each input microwave power, the fitted curve reproduces the experimental data precisely, as shown by the solid black lines in Fig. 5(a).
@@
In Fig. 5(c), we plot the microwave amplitude distribution from spin waves $B_{\mathrm{sw}}(x)$, estimated from the fitting results at eleven different input microwave powers. The left end of the plot is set at $x=0.91\,\mathrm{\mu m}$. Notably, the amplitude at the left end continues to increase up to the maximum input power of $P_{\mathrm{mw}}=22\,\mathrm{dBm}$. In contrast, the amplitude at the right end saturates at around $B_{\mathrm{sw}}=0.5\,\mathrm{mT}$, indicating the onset of nonlinearity. Figure 5(d) shows the estimated wavenumber distribution $k_{x}(x)$ for each input microwave power. The modulation of $k_{x}(x)$ across the field of view is minimal; even at the maximum input microwave power of $P_{\mathrm{mw}}=22\,\mathrm{dBm}$, the variation from the left to the right end is less than $0.1\,\mathrm{rad/\mu m}$. These observations are markedly different from those for YIG-A, as shown in Figs. 2(c) and (d).
@@
Figure 6: (a) Calculated spin-wave amplitude distribution $m_{x}(x)$ for each input microwave power for YIG-B. (b) The heatmap of the spin-wave amplitude distribution for each input microwave amplitude for YIG-B. (c) (d) Dependence of the spin-wave amplitude on input microwave amplitude (c) at $x=4\,\mathrm{\mu m}$ and (d) at $x=78\,\mathrm{\mu m}$ for YIG-B.
@@
As we did for YIG-A, we use Eq. (5) to convert $B_{\mathrm{sw}}(x)$, shown in Fig. 5(c), into the corresponding spin-wave amplitude distributions $m_{x}(x)$. Figure 6(a) shows the converted spin-wave amplitude, and Fig. 6(b) displays a heatmap of $m_{x}(x)$ for all data, together with the individual profiles from Fig. 6(a). To investigate the nonlinearity of the spin waves, we examine the dependence of the spin-wave amplitude on the input microwave amplitude at two positions: $x=4\,\mathrm{\mu m}$ on the left side of the field of view and $x=78\,\mathrm{\mu m}$ on the right side. Figure 6(c) shows the dependence of the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ on the input microwave amplitude. The red dashed line represents a linear fit through the origin, obtained using data from the five smallest input microwave amplitudes (up to $P_{\mathrm{mw}}=10\,\mathrm{dBm}$). The data indicate that nonlinearity begins to occur when the spin-wave amplitude exceeds approximately $10\,\mathrm{mT}$. Although the amplitude does not immediately saturate, it deviates from the linear trend and the gradient continues to decrease. Similarly, Fig. 6(d) shows the dependence of the spin-wave amplitude at $x=78\,\mathrm{\mu m}$. As for the data at $x=4\,\mathrm{\mu m}$, nonlinearity appears when the amplitude reaches around $10\,\mathrm{mT}$. However, in contrast to the left side, the amplitude at $x=78\,\mathrm{\mu m}$ does not increase significantly in the nonlinear region and appears to saturate around $13\,\mathrm{mT}$. These behaviors are qualitatively similar to those observed for YIG-A, but the threshold at which nonlinearity emerges is smaller than in YIG-A.
@@
We compare the obtained results with the theoretical threshold given by Eq. (14). The right column of Table 1 lists the parameters used for the calculation. The saturation magnetization is taken as $M_{\mathrm{s}}=162\,\mathrm{mT}$, as measured by SQUID magnetometry. The anisotropy field is estimated from the difference between the effective magnetization $M_{\mathrm{eff}}=177.6\,\mathrm{mT}$ and the saturation magnetization. The decay rate is calculated from the decay length $l_{d}=318\,\mathrm{\mu m}$, obtained from the fitting result at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$. With these parameters, the threshold spin-wave amplitude for the onset of four-magnon decay is estimated to be $\mu_{0}m^{\mathrm{th}}_{x}=10.5\,\mathrm{mT}$. As expected, this threshold is lower than that of YIG-A, reflecting the smaller decay rate in the present case. As shown in Figs. 6(c) and (d), the experimental data indicate that nonlinearity begins to occur when the spin-wave amplitude exceeds approximately $10\,\mathrm{mT}$ (marked by the horizontal lines in the graphs), again in good agreement with this theoretical prediction.
@@ IV Conclusion and Perspectives
To conclude, we have employed QDM to investigate the nonlinear dynamics of surface spin waves quantitatively. For YIG-A, we swept the input microwave power to map the microwave amplitude distribution, revealing both nonlinear behavior in the amplitude profile and shifts in the wavenumber resulting from a reduction in the saturation magnetization. Quantitative analysis of these data showed that nonlinearity in the amplitude emerges once the spin-wave amplitude exceeds a specific threshold, consistent with theoretical predictions based on four-magnon scattering. Furthermore, as the spin-wave amplitude increases, we observed modulation of the wavenumber within the field of view, and our quantitative analysis revealed that this modulation largely originates from the spin waves generated by multi-magnon scattering. Similar measurements conducted on YIG-B, which exhibits a different relaxation rate, confirmed that the threshold for the onset of nonlinearity also agrees with the theoretical predictions.
@@
These results deepen our understanding of nonlinear spin-wave dynamics and underscore the utility of NV centers as sensors for the quantitative investigation of nonlinear spin-wave phenomena, which have previously been accessible only at a qualitative level.
@@
As a future extension of this work, one could consider the quantitative visualization of the incoherent magnons generated by four-magnon scattering. NV centers can quantitatively measure the magnetic noise at their resonance frequency via longitudinal relaxation measurements. Therefore, by exciting surface spin waves with frequencies detuned from the NV spin resonance and monitoring the relaxation rate of the NV spins while varying the input microwave power, it becomes possible to extract quantitative information about the incoherent magnons produced by nonlinear scattering [37]. Furthermore, under the same conditions, quantitative detection of the coherent spin waves detuned from the NV resonance frequency is possible by employing sensing based on the AC Zeeman effect [26].
@@
In addition, nonlinear phenomena regarding spin waves are not limited to surface spin waves. For example, parametric excitation, where spin waves are generated by microwave fields at twice the frequency of the spin waves oriented along the saturation magnetization direction [3, 4, 5], is an important example. Furthermore, in the case of large-amplitude spin waves, the propagation of spin-wave bullets, which behave as solitons [44], represents another notable nonlinear effect. NV centers are expected to serve as powerful probes for investigating these phenomena.
The nonlinear dynamics of spin waves are also anticipated to have applications in spintronics devices [45, 46]. Recent developments of magnonics has been especially active in spin waves in the shorter-wavelength, higher-frequency regime [47], within which nonlinearity of the spin waves in particular holds substantial potential for efficient spin-wave excitation [48] and applications for logic devices [49] and neuromorphic computing [50]. Methods capable of quantitatively evaluating these operations can provide valuable guidelines for device design.
@@ Acknowledgements
This work was partially supported by JST, CREST Grant No. JPMJCR23I2, Japan; Grants-in-Aid for Scientific Research (Nos. JP22K03524, JP22KJ1058, JP22KJ1059, JP23K25800, JP23K21077, JP24K21194, JP24H01666 and JP25H01248); the Mitsubishi Foundation (Grant No. 202310021); Seiko Instruments; Advanced Technology Foundation Research Grants; the Cooperative Research Project of RIEC, Tohoku University; “Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM)” (No. JPMXP1222UT1131) of the Ministry of Education, Culture, Sports, Science and Technology of Japan (MEXT). K.O. and M.T. acknowledge supports from FoPM, WINGS Program, The University of Tokyo.